OpenStudy (anonymous):

Let g(x)=-x^2+3x+5 find g(

OpenStudy (anonymous):

find g(???

OpenStudy (anonymous):

\[g(\pi)\]

OpenStudy (anonymous):

g(x =- x^2+3x+5 => g (π) = -π^2 +3π + 5

OpenStudy (anonymous):

Whatever it's, just plug it in, pretending x as the holder, holding any replaced value!